A line is tangent to the circle whose equation is x^+y^+6x-6y-2=0 at the point (1,5). Find the Slope-intercept form of the tangent line.
Can you help??

Answer by venugopalramana(3286)

A line is tangent to the circle whose equation is x^+y^+6x-6y-2=0 at the point (1,5). Find the Slope-intercept form of the tangent line.
Can you help?
I THINK YOU MEAN
LET S=x^2+y^2+6x-6y-2=0 at the point (1,5). Find the Slope-intercept form of the tangent line...PLEASE TYPE PROPERLY....ASSUMING SO THE EQN.OF TANGENT AT (X1,Y1)
ON THE CIRCLE IS GIVEN BY S1=0..SO
XX1+YY1+3(X+X1)-3(Y+Y1)-2=0
X+5Y+3(X+1)-3(Y+5)-2=0
4X+2Y-14=0
Y=-2X+7....IS THE EQN. OF TANGENT